0759079776:

Blue light (λ = 475 nm) is sent through a single slit with a width of 2.1 µm. What is the maximum possible number of bright fringes produced on the screen?

SemiDefinite:

For single-slit width 'w', and destructive interference: \(w \sin \theta = m \lambda \qquad m \in \mathbb Z, m \ne 0\) At extreme \(\theta = \frac{\pi}{2}\) \(m = \frac{w}{\lambda} \approx 4.4 \) That makes for 7 bright fringes Given that the central fringe is bright, & ors are arranged symmetrically